the problems is 3333333334 to the 2 power is a 20 digit answer. what is the sum of the digits. i cant get the 20 digit answer any help is greatly appreciated

Answer by stanbon(48535)

See if you can find a pattern in the sum of the digits for:
(3+1)^2 = 3^2+3(3)+1
(33+1)^2= 33^2+2(33)+1
(333+1)^2= 333^2+2(33)+1
etc.
Your number is (3333333333+1)^2
I got an answer of sum of digits is 61 but
I may have made a mistake along the line.
Cheers,
Stan H.